1. A darevevil plans to perform a very dangerous stunt by diving (viz., falling) from a height of 11:9 m (above the waterís surface) into a small pool Ölled with water to a depth of 30 cm. What minimum deceleration will be required (in the water) to advert certain, mortal failure?
2. (a) Determine the average number of gís experienced by someone abruptly coming to rest from an initial velocity of 30 mph (such as in a car crash). Take the distance involved in coming to rest to be 10 cm. (b) Determine the time it takes to stop.
1
Expert's answer
2019-01-25T12:02:24-0500

1.

The speed before collision:

v=√2gH

From the conservation of energy:

v=√2ah

Thus:

a=g H/h=9.8 11.9/0.3=389 m/s^2


2.

a)

a=v^2/2d

a/g=v^2/2gd=(30∙0.44704)^2/(2(9.8)(0.1))=92

b)

t=2d/v=(2(0.1))/((30∙0.44704) )=0.015 s.